OpenStudy (anonymous):

What is the simplified form of m radical 16?

OpenStudy (anonymous):

@asnaseer could u help me?

OpenStudy (asnaseer):

so question is - what is simplified form of: \(m\sqrt{16}\)?

OpenStudy (anonymous):

Yes

OpenStudy (asnaseer):

do you what the square root of 16 is?

OpenStudy (anonymous):

4

OpenStudy (asnaseer):

correct - so what do you think the answer should be?

OpenStudy (anonymous):

Wait , the questione was radical m^16

OpenStudy (anonymous):

I'm sorry I typed it wrong

OpenStudy (asnaseer):

so: \(\sqrt{m^{16}}\)?

OpenStudy (anonymous):

Yessir

OpenStudy (asnaseer):

ok - do you know this rule for indices?\[a^m\times a^n=a^{m+n}\]

OpenStudy (anonymous):

I kinda get it.. Confuses me though a little

OpenStudy (asnaseer):

think of it like this:\[a^3\times a^4=(a\times a\times a)\times(a\times a\times a\times a)\]\[\qquad=a\times a\times a\times a\times a\times a\times a\]\[\qquad=a^7=a^{3+4}\]

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

Oh yes

OpenStudy (asnaseer):

ok, now a square root of X is a number which, when it multiplied by itself, will give you back X. agreed?

OpenStudy (anonymous):

Ok yes..

OpenStudy (asnaseer):

so we need to find some power of m, which, when it is multiplied by itself, will give you \(m^{16}\), i.e. we need to find x such that:\[m^x\times m^x=m^{16}\]then \(m^x\) would be the square root of \(m^{16}\) - agreed?

OpenStudy (anonymous):

Yea.. So 4*4 is 16' so so is 6 andn2 and 1 and 16 and 8 and 2

OpenStudy (asnaseer):

?

OpenStudy (asnaseer):

\[m^x\times m^x=m^{x+x}=m^{2x}\]

OpenStudy (anonymous):

4&4 multiply to be 16 so does 8&2 so does 6&2

OpenStudy (asnaseer):

we ADD powers - look at the examples we did above

OpenStudy (asnaseer):

remember: \(a^m\times a^n=a^{m+n}\)

OpenStudy (anonymous):

Oolong so 4+4 = 8 and 8* 2 is 16

OpenStudy (asnaseer):

why are you multiplying?

OpenStudy (anonymous):

Oh so it's just m^8

OpenStudy (asnaseer):

we need to find x such that:\[m^x\times m^x=m^{16}\]i.e.:\[m^{2x}=m^{16}\]which implies:\[2x=16\]

OpenStudy (asnaseer):

yes - thats right

OpenStudy (anonymous):

Oh ok cool

OpenStudy (asnaseer):

this also highlights that square root of x is just x to the power of a half, i.e.:\[\sqrt{x}=x^{\frac{1}{2}}\]

OpenStudy (asnaseer):

we could have used this approach too

OpenStudy (asnaseer):

this would require the following rule:\[(a^m)^n=a^{mn}\]

OpenStudy (anonymous):

Oh I see... Makes sense could u help me again if ur not busy? :)

OpenStudy (anonymous):

I like that way too!

OpenStudy (asnaseer):

so we get:\[\sqrt{m^{16}}=(m^{16})^{\frac{1}{2}}=m^{16\times\frac{1}{2}}=m^8\]

OpenStudy (anonymous):

Easier ! Thanks so much

OpenStudy (asnaseer):

yw :) ok - just post your new question in the list to the left. I need to check a few things on the site - so if no one has helped you by the time I am back then I'll come to help you.